OpenStudy (anonymous):

okay guys please answer this for me: the length of a rectangle is 6 centimeters more than the width. the area is 216 square centimerters. what are the dementions??? oh and if you could work it out for me too thatd be great!!!!

OpenStudy (anonymous):

L = W + 6 LW = 216 (W+6)w = 216 w^2 + 6w = 216 w^2 + 6w - 216 = 0 (W - 12) (W + 18) = 0 w = 12 or -18 -18 makes no sense, so W = 12 L = 12 + 6 L = 18

OpenStudy (anonymous):

thanks so much that actually looks right :)